y=(u-2)3 and u=1/2x=1


1
Expert's answer
2021-06-25T07:22:23-0400

"Given,y=(u-2)^3.\\
\text{Substitute}\space u=\frac{1}{2x}=1\\\space \text{in the given equation, we obtain,}\\
y=(\frac{1}{2x}-2)^3\\
y=(1-2)^3 \space( given\frac{1}{2x}=1)\\
y=(-1)^3\\
y=-1"